# 11affiliationtext: IMT School for Advanced Studies Lucca22affiliationtext: Universidad de Alicante

### v1

@@ 3 Simulation Study
P1 0.706 0.630 0.958 0.916 0.856 1.000 0.996 0.978 1.000
P2 0.544 0.442 0.802 0.720 0.676 0.982 0.942 0.886 1.000
P3 0.430 0.238 0.348 0.716 0.384 0.388 0.986 0.644 0.392
P4 0.684 0.426 0.530 0.972 0.628 0.518 1.000 0.950 0.508
Table 1 shows evidence that the M-vine test and the linear test control the empirical size under or around the significance level of $\alpha=0.05$ better than the S-vine test. In terms of power, the M-vine test outperforms the S-vine test, in every power assessment model, in small and large samples, with the biggest difference when working with series from the data-generating model P3, that exhibit a quadratic dependence relationship. Furthermore, the power of the M-vine test seems to increase as the sample size grows faster than the one of the S-vine test. Notice that the M-vine and the S-vine tests have the same bootstrapped version of the test statistic, i.e. $\widehat{GC}^{0\;mean}_{Y\rightarrow X}$, and so the difference in the performance relies upon the computation of $\widehat{GC}^{mean}_{Y\rightarrow X}$, that explains the difference in the obtained $p$-values. When compared to the Linear test, the M-vine test tends to have a smaller power across the assessment models P1 and P2 that corresponds to linear or roughly linear models, respectively; this difference is decreased as we move to larger samples, becoming negligible at $T=200$. On the contrary, the power of M-vine test is substantially bigger than the one of the Linear test when working with non-linear model specifications, such as P3 and P4, particularly when working with larger samples: indeed, the power of the Linear test does not increase as the sample size grows, remaining considerably low. Hence, the presented analyses show that the M-vine test is an excellent tool for detecting the Granger causality in the presence of non-linear dependencies, filling in a gap present in the current literature.
@@
rowodd = c, row4 = c, cell11 = r=2, cell12 = c=3, cell15 = c=3, cell18 = c=3, cell28 = c, cell29 = c, cell210 = c, vline5,8 = 2-5, hline1,6 = -, hline2-3 = 2-10, Model & T=50 T=100 T=200
@@
rowodd = c, row4 = c, row6 = c, cell11 = r=2, cell12 = c=3, cell15 = c=3, cell18 = c=3, vline5,8 = 2-6, hline1,7 = -, hline2-3 = 2-10, Model & T=50 T=100 T=200
@@
Table 3 and 4 show the average $p$-value across simulations for each size and power assessment model, respectively, together with their empirical standard deviation. One can see that, on average, the $p$-values for all methods are around the significance level $\alpha=0.05$ in every size assessment model. However, those of the M-vine approach have the lower standard deviation among the three tests in all size models, both in small and larger samples. On the other hand, from Table 4, we can see that for the linear (P1) and close to linear (P2) model specifications, the $p$-values of the Linear method are the ones that are, on average, further below the predefined significance level, also displaying the lower standard deviation. When working with non-linear dependence structures (P3 and P4), this no longer holds as the $p$-values for the M-vine procedure are, on average, closer to the significance level, and they exhibit the lower empirical deviation out of the three methods. In Figure 1 we plot the distributions of the $p$-values obtained with each method in cases P3 and P4 with $T=100$ in order to further support the previous considerations.
@@
Thus far, we have assumed that $(X_{t},Y_{t})_{t}$ is a first-order Markov process (i.e. $k=1$), and, under this condition, the M-vine test results to have good statistical properties for sample size $T\geq 100$. In order to analyse how these properties change when working with higher-order Markov processes, we repeat the simulation study for the case $k=4$. The selection of this particular Markov order is driven by the empirical application from Section 4, as it is the only other Markov order, besides $k=1$, obtained by our selection procedure when fitting the M-vine models to the U.S. macroeconomic data (we refer to Section 4.1 for a detailed description of the data). To adjust the simulation study for $k=4$, we modified the previously presented size and power assessment models so that they represent the same dependence structure scenarios, but letting $(X_{t},Y_{t})_{t}$ be a bivariate Markov stationary stochastic process of order $k=4$. The modified assessment models are given in Appendix A.2. Table 5 and Table 6 show the empirical size and power of each test when applied to a bivariate Markov stationary process of order $k=4$. From them, we can see that overall the results exhibit the same pattern as the case of $k=1$, with the M-vine test still outperforming in terms of power the S-vine test in both linear and non-linear assessment models. Moreover, when compared to the linear test, the M-vine test has a comparable performance when the processes have a linear or almost linear dependence and the sample size is large enough, while it performs better in non-linear cases. However, we note that for the order $k=4$, the power of the M-vine test seems to converge to 1 at a slower pace than in the case of $k=1$. Thus, for $k=4$, we require a greater $T$ to let the M-vine test have a good power in the cases P3 and P4.
@@ A.1 Vine copula models
Copula models constitute a type of dependence-capturing models, which in contrast to VAR allow for non-linear dependences when working with high dimensional data. Their advantage relies on the fact that the univariate distributions of the elements from a multivariate system can be selected freely and then model the dependence structure independently of these marginals by linking them through an appropriate copula function. However, copula functions have their own limitations. Particularly, when dealing with high-dimensional random vectors the set of available parametric copulas is considerably limited and so the selection of a suitable high-dimensional copula function could result difficult. Vine copulas come as an alternative: since the large amount of existing bivariate copula families to be employed, the vine copula approach consists in the decomposition of a multivariate copula density into a product of bivariate copula densities through the use of the pair-copula construction (Joe (1996)). Nevertheless, the number of possible decompositions becomes difficult to track and organize: in fact, the number of possible decompositions increases super-exponentially with the dimension of the random vector (Morales Napoles, 2010). Then, Regular vines (R-vines) were introduced as a graphical way of specifying the sequence of trees that characterises all valid factorizations of a $d$-dimensional copula. Briefly, as defined on Cooke and Kurowicka (2006) an R-vine on $d$ random variables is a set of graphs $\mathcal{V}=(T_{1};\dots,T_{d-1})$ where each tree $T_{i}$ has a node set $N_{i}$ and an edge set $E_{i}$, such that:
@@ A.1.2 A sub-class of S-vine copulas: M-vine copulas
The main difference between M-vine and S-vine copulas rely on the fact that the tree structures for a M-vine are fixed as the cross-sectional vines have to be connected across time periods through an edge between two nodes lying in the same end of the vine. On the other hand, S-vines allow for general cross-sectional structures (not necessarily D-vines) and provide a higher degree of freedom for connecting them across time compared to M-vines. This indeed provide more flexibility for S-vines in order to fit different kind of data; however, in cases when we need to guarantee the presence of one or some specific copulas in the vine structure, this can be an undesirable feature: for instance, in the bootstrap procedure from Part B of Section 2.1, where the copulas $c_{X_{t},X_{t+1}}$ and $c_{X_{t},Y_{t}}$ are needed to generate bootstrapped samples under the null hypothesis, using an S-vine structure is not feasible as these two copulas might not appear in its first tree. Therefore, even if S-vines constitute a more general class of vine copulas that can ensure stationarity of a time series (in which M-vines are a sub-class), for some cases where we need to impose some particular conditions on the dependence structure, fitting a more restrictive vine structure (e.g. an M-vine) could be beneficial.
A.2 Size and power assessment models for $k=4$
@@
A.3 Previous literature: Granger-causality relationship between Energy Consumption and Economic Growth
@@
A.4 Preliminary time series analysis for the application

### v2

@@ 3 Simulation Study
row1 = c, column4 = c, column8 = c, column12 = c, cell11 = r=2, cell12 = c=4, cell16 = c=4, cell110 = c=4, cell32 = c, cell33 = c, cell35 = c, cell36 = c, cell37 = c, cell39 = c, cell310 = c, cell311 = c, cell313 = c, cell42 = c, cell43 = c, cell45 = c, cell46 = c, cell47 = c, cell49 = c, cell410 = c, cell411 = c, cell413 = c, cell52 = c, cell53 = c, cell55 = c, cell56 = c, cell57 = c, cell59 = c, cell510 = c, cell511 = c, cell513 = c, cell62 = c, cell63 = c, cell65 = c, cell66 = c, cell67 = c, cell69 = c, cell610 = c, cell611 = c, cell613 = c, cell72 = c, cell73 = c, cell75 = c, cell76 = c, cell77 = c, cell79 = c, cell710 = c, cell711 = c, cell713 = c, vline3,7,11 = 1, vline6,10 = 2-7, hline1,8 = -, hline2-3 = 2-13, Model & T=50 T=100 T=200
M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear
P1 0.706 0.630 0.000 0.958 0.916 0.856 0.038 1.000 0.996 0.978 0.172 1.000
P2 0.544 0.442 0.002 0.802 0.720 0.676 0.090 0.982 0.942 0.886 0.378 1.000
P3 0.430 0.238 0.000 0.348 0.716 0.384 0.034 0.388 0.986 0.644 0.172 0.392
P4 0.684 0.426 0.000 0.530 0.972 0.628 0.176 0.518 1.000 0.950 0.612 0.508
P5 0.142 0.084 0.002 0.092 0.198 0.146 0.024 0.064 0.314 0.214 0.128 0.052
Table 1 shows evidence that the M-vine test and the linear test control the empirical size below or around the significance level of $\alpha=0.05$ better than the S-vine test. Furthermore, the NlinTS test is the one that exhibits higher size distortions; in fact, when compared to the other tests, only the S-vine test by Jang et al. (2022) exhibits in some cases a size greater than or equal to $0.070$. In terms of power, the NlinTS test is the one with the worst performance. Moreover, the M-vine test outperforms the S-vine test, in every power assessment model, in small and large samples, with the biggest difference when working with series from the data-generating model P3, that exhibit a quadratic dependence relationship. Furthermore, the power of the M-vine test seems to increase as the sample size grows faster than the one of the S-vine test. Notice that the M-vine and the S-vine tests have the same value of the test statistic under the null hypothesis, i.e. $\widehat{GC}^{0\;mean}_{Y\rightarrow X}$, and so the difference in the performance relies upon the computation of $\widehat{GC}^{mean}_{Y\rightarrow X}$, that explains the difference in the obtained $p$-values. When compared to the linear test, the M-vine test tends to have a smaller power across the assessment models P1 and P2 that correspond, respectively, to a linear model and to a non-linear model with a bounded non-linear term; this difference is decreased as we move to larger samples, becoming negligible at $T=200$. On the contrary, the power of M-vine test is substantially bigger than the one of the linear test when working with non-linear model specifications (precisely, with models exhibiting a quadratic dependence), such as P3, P4, and P5, particularly when working with larger samples: indeed, the power of the linear test does not increase as the sample size grows, remaining considerably low. Notably, all the tests considered in the analysis exhibit low power working with the P5 model, in which the weight of the term $Y_{t-1}$ in the dynamics of the process $X$ is lower than the one in the previous models, which implies that $Y$ Granger-causes $X$ to a lesser extent, requiring a larger sample size to achieve a higher power. Hence, the presented analyses show that the M-vine test brings remarkable improvements in the current literature regarding tools for detecting the Granger causality in the presence of non-linear dependencies.
@@
rowodd = c, row4 = c, row6 = c, cell11 = r=2, cell12 = c=4, cell16 = c=4, cell110 = c=4, cell210 = c, cell211 = c, cell213 = c, vline6,10 = 2-7, hline1,8 = -, hline2-3 = 2-13, Model & T=50 T=100 T=200
M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear
@@
rowodd = c, row4 = c, row6 = c, cell11 = r=2, cell12 = c=4, cell16 = c=4, cell110 = c=4, cell24 = c, cell28 = c, cell212 = c, vline6,10 = 2-7, hline1,8 = -, hline2-3 = 2-13, Model & T=50 T=100 T=200
M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear
@@
Table 3 and 4 show the average $p$-value across simulations for each size and power assessment model, respectively, together with their empirical standard deviation. One can see that, on average, the $p$-values for all methods are around the significance level $\alpha=0.05$ in every size assessment model (with the exception of P5 due to the sample size as previously mentioned). However, those of the M-vine approach have the lower standard deviation among the three tests in all size models, both in small and larger samples. On the other hand, from Table 4, we can see that for the linear (P1) and close to linear (P2) model specifications, the $p$-values of the Linear method are the ones that are, on average, further below the predefined significance level, also displaying the lower standard deviation. When working with non-linear dependence structures (P3, P4 and P5), this no longer holds as the $p$-values for the M-vine procedure are, on average, closer to the significance level, and they exhibit the lower empirical deviation out of the three methods. In Figure 1 we plot the distributions of the $p$-values obtained with each method in cases P3 and P4 with $T=100$ in order to further support the previous considerations.
@@
Thus far, we have assumed that $(X_{t},Y_{t})_{t}$ is a first-order Markov process (i.e. $k=1$), and, under this condition, the M-vine test results to have good statistical properties for sample size $T\geq 100$. In order to analyse how these properties change when working with higher-order Markov processes, we repeat the simulation study for the case $k=4$. The selection of this particular Markov order is driven by the empirical application from Section 4, as it is the only other Markov order, besides $k=1$, obtained by our selection procedure when fitting the M-vine models to the U.S. macroeconomic data (we refer to Section 4.1 for a detailed description of the data). To adjust the simulation study for $k=4$, we modified a selection of the previously presented size and power assessment models so that they represent the same dependence structure scenarios, but letting $(X_{t},Y_{t})_{t}$ be a bivariate Markov stationary stochastic process of order $k=4$. The modified assessment models are given in Appendix A.3. Table 5 and Table 6 show the empirical size and power of each test when applied to a bivariate Markov stationary process of order $k=4$. From them, we can see that overall the results exhibit the same pattern as the case of $k=1$, with the M-vine test still outperforming in terms of power the S-vine test in both linear and non-linear assessment models. Moreover, when compared to the linear test, the M-vine test has a comparable performance when the processes have a linear or almost linear dependence and the sample size is large enough, while it performs better in non-linear cases. However, we note that for the order $k=4$, the power of the M-vine test seems to converge to 1 at a slower pace than in the case of $k=1$. Thus, for $k=4$, we require a greater $T$ to let the M-vine test have a good power in the cases P3 and P4.
@@ A.1 Vine copula models
Copula functions are the mathematical tool that describes any type of dependence structure. Hence, in contrast to VAR, copula models allow for non-linear dependences when working with high dimensional data.
Moreover, Sklar’s theorem (Sklar, 1959) shows that the marginal distributions of the random variables in a multivariate system can be selected freely and then the dependence structure among these marginals can be modeled by linking them through an appropriate copula function.
Theorem 1 (Sklar’s Theorem).
Consider a $d$-dimensional joint cumulative distribution function $H$ with one-dimensional marginals $F_{1},\dots,F_{d}$. Then there exists a d-copula $C$ such that for $(x_{1},\dots,x_{d})\in\mathbb{R}^{d}$,
$H(x_{1},\dots,x_{d})=C(F_{1}(x_{1}),\dots,F_{d}(x_{d})).$
On the other hand, the converse is also true, i.e. consider a copula $C$ and univariate cumulative distribution functions $F_{1},\dots,F_{d}$. Then $H$ is a multivariate cumulative distribution function with marginals $F_{1},\dots,F_{d}$.
However, copula functions have their own limitations when we have to select and estimate them by means of the data. Particularly, when dealing with high-dimensional random vectors the set of available parametric copulas is considerably limited and so the selection of a suitable high-dimensional copula function could result difficult. Vine copulas come as an alternative: since the large amount of existing bivariate copula families to be employed, the vine copula approach consists in the decomposition of a multivariate copula density into a product of bivariate copula densities through the use of the pair-copula construction (Joe (1996)). Nevertheless, the number of possible decompositions becomes difficult to track and organize: in fact, the number of possible decompositions increases super-exponentially with the dimension of the random vector (Morales Napoles, 2010). Then, Regular vines (R-vines) were introduced as a graphical way of specifying the sequence of trees that characterises all valid factorizations of a $d$-dimensional copula. Briefly, as defined on Cooke and Kurowicka (2006) an R-vine on $d$ random variables is a set of graphs $\mathcal{V}=(T_{1};\dots,T_{d-1})$ where each tree $T_{i}$ has a node set $N_{i}$ and an edge set $E_{i}$, such that:
@@ A.1.2 A sub-class of S-vine copulas: M-vine copulas
The main difference between M-vine and S-vine copulas relies on the fact that the tree structures for a M-vine are fixed as the cross-sectional vines have to be connected across time periods through an edge between two nodes lying in the same end of the vine. On the other hand, S-vines allow for general cross-sectional structures (not necessarily D-vines) and provide a higher degree of freedom for connecting them across time compared to M-vines. This indeed provide more flexibility for S-vines in order to fit different kind of data; however, in cases when we need to guarantee the presence of one or some specific copulas in the vine structure, this can be an undesirable feature: for instance, in Part B of our statistical procedure illustrated in Section 2.1, where the copulas $c_{X_{t},X_{t+1}}$ and $c_{X_{t},Y_{t}}$ are needed to generate independent samples under the null hypothesis, using an S-vine structure is not feasible as these two copulas might not appear in its first tree. Therefore, even if S-vines constitute a general class of vine copulas that can ensure stationarity of a time series, in which M-vines are a sub-class, in some cases when we need to detect particular dependencies, fitting a more restrictive vine structure (e.g. an M-vine) could be beneficial.
A.1.3 Selection and estimation for M-vine copulas
In order to select the random variables associated to each node of the vine structure, and the pair copula families associated to each edge in the vine structure, one could rely on a preliminary analysis of the data being used. In practice, one could select which variables are associated to each node in the first tree considering the scope of the particular application considered, and then select the copula families associated to each edge using scatter plots of pair of empirical probability integral transforms of the corresponding variables and their lagged values (Beare and Seo, 2015).
Alternatively, one could use the algorithm from Dißmann et al. (2013) in order to select and estimate a vine copula specification, which first fits the strongest dependence based on the absolute value of the empirical Kendall’s $\tau$ between any pair of random variables in the system and, subsequently, the selection of the pair copulas and the estimation of their corresponding parameters is done sequentially through step-wise maximum likelihood estimation (MLE). Firstly, the empirical distribution functions are used to estimate the marginals by
$\widehat{F}_{j}(x_{j})=\frac{1}{T+1}\sum_{t=1}^{T}\mathbb{1}(X_{t,j}\leq x),\qquad t=1,\dots,T,j=1,\dots,d.$
Using these estimates, one can obtain "pseudo-observations" $\widehat{U}=\widehat{F}_{j}(X_{j})$, that are used for the estimation of the copula parameters. The step-wise MLE estimates the parameters of the pair-copulas of each tree $T_{k}$ separately, beginning with the first tree. For the setting of a $k$-Markov process, Nagler et al. (2022) showed that this can be done through
$\widehat{\boldsymbol{\theta}}_{[e^{\prime}]}=\arg\max_{\boldsymbol{\theta}_{\mathbf{[e^{\prime}]}}}\sum_{e\sim e^{\prime}}\log c_{[e]}\left\{C_{a_{[e]}|D_{[e]}}\left(\widehat{U}_{a_{e}}|\mathbf{\widehat{U}}_{D_{e}};\widehat{\boldsymbol{\theta}}_{S_{a}([e])},\right),C_{b_{[e]}|D_{[e]}}\left(\widehat{U}_{b_{e}}|\mathbf{\widehat{U}}_{D_{e}};\widehat{\boldsymbol{\theta}}_{S_{b}([e])},\right);\widehat{\boldsymbol{\theta}}_{[e^{\prime}]}\right\},$
where $\boldsymbol{\theta}_{[e]}$ is the parameter of the copula associated to the edge $e$, $[e]=\{e^{\prime}\sim e\}$ is the equivalence class of edge $e$, containing all the the edges that share the same copula and parameters than $e$. Lastly, $\widehat{\boldsymbol{\theta}}_{S_{a}([e])}$ and $\widehat{\boldsymbol{\theta}}_{S_{b}([e])}$ contain previously estimated parameters from prior trees.
[…]
A.2 Linear bivariate Granger causality test
Let $X=(X_{t})_{t}$, $Y=(Y_{t})_{t}$ be two stationary time series with $t=1,\dots,T$. To test whether $Y$ Granger-causes $X$, the linear test assumes a particular autoregressive lag length $p$, that is
$X_{t}=c_{1}+\sum_{j=1}^{p}\alpha_{j}X_{t-j}+\sum_{j=1}^{p}\beta_{j}Y_{t-j}+u_{t}\,,$
where $u_{t}$ is a white noise, and use estimation by Ordinary Least Squares (OLS). Then, an asymptotic $F$ test of the null hypothesis
$H_{0}:\beta_{1}=\dots=\beta_{p}=0$
is conducted. To implement this test, one needs to compute the sum of squared residuals
$RSS_{1}=\sum_{j=1}^{T}\hat{u}_{t}$
and compare it with the sum of squared residuals of the univariate autoregressive model of order $p$ for $X$,
$RSS_{0}=\sum_{j=1}^{T}\hat{e}_{t}$
where
$X_{t}=c_{0}+\sum_{j=1}^{p}\gamma_{j}X_{t-j}+e_{t}$
is also estimated by OLS. Finally, if
$S=\frac{T(RSS_{0}-RSS_{1})}{RSS_{1}}$
is greater than the 5% critical value for a $\mathcal{X}^{2}(p)$ distribution, then we reject the null hypothesis that Y does not Granger-cause X.
A.3 Size and power assessment models for $k=4$
@@
A.4 Previous literature: Granger-causality relationship between Energy Consumption and Economic Growth
@@
A.5 Preliminary time series analysis for the application